OpenStudy (anonymous):

In laws of motion...if M1,M2,F1 are given then how to find F2...please reply asap

OpenStudy (primeralph):

Need a diagram, angles or the full question.

OpenStudy (anonymous):

If M1 and M2 have same acceleration a then a=F1/M1 = F2/M2 so F2 = F1* M2/M1

OpenStudy (anonymous):

i believe question is incomplete...